# On the Mode-Seeking Properties of Langevin Dynamics

### v1

@@ 4.2 Langevin Dynamics in Sub-Gaussian Mixtures
We further generalize our results to sub-Gaussian mixtures. We impose the following assumptions on the mixture. It is worth noting that these assumptions automatically hold for Gaussian mixtures.
@@
We validate the feasibility of Assumption 2.v. in Lemma 9 in the Appendix. With Assumption 2, we show the mode-seeking tendency of Langevin dynamics under sub-Gaussian distributions in Theorem 3 and defer the proof to Appendix A.3.
@@
Finally, we slightly modify Assumption 2 and extend our results to annealed Langevin dynamics under sub-Gaussian mixtures in Theorem 4. The details of Assumption 3 and the proof of Theorem 4 are deferred to Appendix A.4.
Theorem 4.
@@ 5 Chained Langevin Dynamics
To reduce the mode-seeking tendencies of vanilla and annealed Langevin dynamics, we propose Chained Langevin Dynamics (Chained-LD) in Algorithm 1. While vanilla and annealed Langevin dynamics apply gradient updates to all coordinates of the sample in every step, we decompose the sample into patches of constant size and generate each patch sequentially to alleviate the exponential dependency on the dimensionality. More precisely, we divide a sample ${\mathbf{x}}$ into $d/Q$ patches ${\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(d/Q)}$ of some constant size $Q$, and apply annealed Langevin dynamics to sample each patch ${\mathbf{x}}^{(q)}$ (for $q\in[d/Q]$) from the conditional distribution $P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$.
@@
1:Patch size $Q$, dimension $d$, conditional score function estimator ${\mathbf{s}}_{\bm{\theta}}$, number of iterations $T$, noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$, step size $\left\{{\delta_{t}}\right\}_{t\in[TQ/d]}$.
@@
5:         ${\mathbf{x}}_{t}^{(q)}\leftarrow{\mathbf{x}}_{t-1}^{(q)}+\frac{\delta_{t}}{2}{\mathbf{s}}_{\bm{\theta}}\left({\mathbf{x}}_{t}^{(q)}\mid{\sigma}_{t},{\mathbf{x}}_{t}^{(1)},\cdots,{\mathbf{x}}_{t}^{(q-1)}\right)+\sqrt{\delta_{t}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{Q},{\bm{I}}_{Q})$
@@
7:     ${\mathbf{x}}_{0}^{(q)}\leftarrow{\mathbf{x}}_{TQ/d}^{(q)}$
@@
9:return ${\mathbf{x}}_{TQ/d}$
An ideal conditional score function estimator ${\mathbf{s}}_{\bm{\theta}}$ could jointly estimate the scores of all perturbed conditional patch distribution, i.e., $\forall{\sigma}\in\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]},q\in[d/Q]$,
Then, combining the objectives gives the following loss
As shown in [44], an estimator ${\mathbf{s}}_{\bm{\theta}}$ with enough capacity minimizes the loss ${\mathcal{L}}$ if and only if ${\mathbf{s}}_{\bm{\theta}}$ outputs the scores of all perturbed conditional patch distribution almost surely. Ideally, if a sampler perfectly generates every patch, combining all patches gives a sample from the original distribution since $P({\mathbf{x}})=\prod_{q\in[d/Q]}P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$. In Theorem 5 we give a linear reduction from producing samples of dimension $d$ using Chained-LD to learning the distribution of a $Q$-dimensional variable for constant $Q$. The proof of Theorem 5 is deferred to Appendix A.5.
Theorem 5.
@@
in $\tau(\varepsilon,d)$ iterations for some $\varepsilon>0$. Then, equipped with the sampler algorithm, the Chained-LD algorithm in $\frac{d}{Q}\cdot\tau(\varepsilon,d)$ iterations can achieve
@@ 6 Numerical Results
Figure 1: Samples from a mixture of three Gaussian modes generated by vanilla, annealed, and chained Langevin dynamics. Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 0.
In this section, we empirically evaluated the mode-seeking tendencies of vanilla, annealed, and chained Langevin dynamics. We performed numerical experiments on synthetic Gaussian mixture models and real image datasets including MNIST [45] and Fashion-MNIST [46]. Details on the experiment setup are deferred to Appendix B.
Synthetic Gaussian mixture model: We define the data distribution $P$ as a mixture of three Gaussian components in dimension $d=100$, where mode 0 defined as $P^{(0)}={\mathcal{N}}(\bm{0}_{d},3{\bm{I}}_{d})$ is the universal mode with the largest variance, and mode 1 and mode 2 are respectively defined as $P^{(1)}={\mathcal{N}}(\bm{1}_{d},{\bm{I}}_{d})$ and $P^{(2)}={\mathcal{N}}(-\bm{1}_{d},{\bm{I}}_{d})$. The frequencies of the three modes are 0.2, 0.4 and 0.4, i.e.,
@@
As shown in Figure 1, vanilla and annealed Langevin dynamics cannot find mode 1 or 2 within $10^{6}$ iterations if the sample is initialized in mode 0, while chained Langevin dynamics can find the other two modes in 1000 steps and correctly recover their frequencies as gradually increasing the number of iterations. In Appendix B.1 we present additional experiments on samples initialized in mode 1 or 2, which also verify the mode-seeking tendencies of vanilla and annealed Langevin dynamics.
@@
Image datasets: We construct the distribution as a mixture of two modes by using the original images from MNIST/Fashion-MNIST training dataset (black background and white digits/objects) as the first mode and constructing the second mode by i.i.d. randomly flipping an image (white background and black digits/objects) with probability 0.5. Regarding the neural network architecture of the score function estimator, for vanilla and annealed Langevin dynamics we use U-Net [47] following from [2]. For chained Langevin dynamics, we proposed to use Recurrent Neural Network (RNN) architectures. We note that for a sequence of inputs, the output of RNN from the previous step is fed as input to the current step. Therefore, in the scenario of chained Langevin dynamics, the hidden state of RNN contains information about the previous patches and allows the network to estimate the conditional score function $\nabla_{{\mathbf{x}}^{(q)}}\log P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$. More implementation details are deferred to Appendix B.2.
@@
Figure 2: Samples from a mixture distribution of the original and flipped images from the MNIST dataset generated by vanilla, annealed, and chained Langevin dynamics. The samples are initialized as original images from MNIST.
The numerical results on image datasets are shown in Figures 2 and 3. Vanilla Langevin dynamics fails to generate reasonable samples, as also observed in [2]. When the sample is initialized as original images from the datasets, annealed Langevin dynamics tends to generate samples from the same mode, while chained Langevin dynamics can generate samples from both modes. Additional experiments are deferred to Appendix B.2.
@@
Figure 3: Samples from a mixture distribution of the original and flipped images from the Fashion-MNIST dataset generated by vanilla, annealed, and chained Langevin dynamics. The samples are initialized as original images from Fashion-MNIST.
@@ 7 Conclusion
In this work, we theoretically and numerically studied the mode-seeking properties of vanilla and annealed Langevin dynamics sampling methods under a multi-modal distribution. We characterized Gaussian and sub-Gaussian mixture models under which Langevin dynamics are unlikely to find all the components within a sub-exponential number of iterations. To reduce the mode-seeking tendency of vanilla Langevin dynamics, we proposed Chained Langevin Dynamics (Chained-LD) and analyzed its convergence behavior. Studying the connections between Chained-LD and denoising diffusion models will be an interesting topic for future exploration. Our RNN-based implementation of Chained-LD is currently limited to image data generation tasks. An interesting future direction is to extend the application of Chained-LD to other domains such as audio and text data. Another future direction could be to study the convergence of Chained-LD under an imperfect score estimation which we did not address in our analysis.
@@ Appendix A Theoretical Analysis on the Mode-Seeking Tendency of Langevin Dynamics
We begin by introducing some well-established lemmas used in our proof. We first provide the proof of Proposition 1 for completeness:
Proof of Proposition 1.
By the definition in equation 2, we have
We use the following lemma on the tail bound for multivariate Gaussian random variables.
@@ A.1 Proof of Theorem 1: Langevin Dynamics under Gaussian Mixtures
We start by proving that the initialization of the state ${\mathbf{x}}_{0}$ has a large norm on the null space with high probability in the following proposition.
@@
A.2 Proof of Theorem 2: Annealed Langevin Dynamics under Gaussian Mixtures
Similar to the proof of Theorem 1, we decompose
Proposition 4.
Proof of Proposition 4.
    we have that for all $t\in[T]$,
    By Lemma 8 we have that with probability at least $1-\exp(-\Omega(d))$,
    Combining the two inequalities implies the desired bound
Therefore we complete the proof of Proposition 4. ∎

### v2

@@ 4.2 Langevin Dynamics in Sub-Gaussian Mixtures
We further generalize our results to sub-Gaussian mixtures. We impose the following assumptions on the mixture. It is worth noting that Assumptions 2.i.-iii. automatically hold for Gaussian mixtures such that $P^{(i)}={\mathcal{N}}({\bm{\mu}}_{i},{\nu}_{i}^{2}I)$, and Assumptions 2.iv. and v. are specific assumptions on the mean and variance of $P^{(i)}$, similar to Assumption 1.
@@
The feasibility of Assumption 2.v. is validated by Lemma 9 in Appendix A.3. With Assumption 2, we show the mode-seeking tendency of Langevin dynamics under sub-Gaussian distributions in Theorem 2 and defer the proof to Appendix A.3.
@@
We remark that an implication of Theorem 2 is the potential difficulty of transition between low-variance modes for Langevin dynamics. For instance, suppose $P^{(1)},\cdots,P^{(k)}$ have bounded support sets with small radius and sufficiently distant means. If the sample is initialized in a low-variance mode $P^{(m)}$ (for $m\in[k]$), either it stays in $P^{(m)}$ and cannot capture other modes, or it escapes $P^{(m)}$ (due to the random noise $\bm{\epsilon}$) and is expected to need to explore the whole space until finding the support sets of the other bounded modes.
Furthermore, in Appendix B we extend our theoretical analysis to annealed Langevin dynamics with bounded noise levels, indicating the effect of annealing noise levels on the mode-seeking tendencies of Langevin dynamics. Aligning with the empirical analysis in [26], we show that bounded noise levels will have a limited impact on Langevin dynamics since they exhibit similar mode-seeking tendencies. On the other hand, as suggested by [26], annealed Langevin dynamics with a significantly larger initial noise level could capture more modes, which is also confirmed by our numerical results in Section 6.
@@ 5 Chained Langevin Dynamics
To reduce the mode-seeking tendencies of vanilla Langevin dynamics, we propose Chained Langevin Dynamics (Chained-LD) in Algorithm 1. While vanilla Langevin dynamics apply gradient updates to all coordinates of the sample in every step, we decompose the sample into patches of constant size and generate each patch sequentially to alleviate the exponential dependency on the dimensionality. More precisely, we divide a sample ${\mathbf{x}}$ into $d/Q$ patches ${\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(d/Q)}$ of some constant size $Q$, and apply Langevin dynamics to sample each patch ${\mathbf{x}}^{(q)}$ (for $q\in[d/Q]$) from the conditional distribution $P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$. Intuitively, vanilla Langevin dynamics needs to explore the entire space (of volume exponentially large in $d$) to find the missing modes, while Chained-LD could significantly lower the volume by dimensionality reduction.
@@
1:Patch size $Q$, dimension $d$, conditional score function $\nabla\log P_{{\sigma}_{t}}$, number of iterations $T$, noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$, step size $\left\{{\delta_{t}}\right\}_{t\in[TQ/d]}$.
@@
5:         ${\mathbf{x}}_{t}^{(q)}\leftarrow{\mathbf{x}}_{t-1}^{(q)}+\frac{\delta_{t}}{2}\nabla\log P_{{\sigma}_{t}}\left({\mathbf{x}}_{t-1}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}\right)+\sqrt{\delta_{t}}\bm{\epsilon}_{t}$, where $\bm{\epsilon}_{t}\sim{\mathcal{N}}(\bm{0}_{Q},{\bm{I}}_{Q})$
@@
7:     ${\mathbf{x}}^{(q)}\leftarrow{\mathbf{x}}_{TQ/d}^{(q)}$
@@
9:return ${\mathbf{x}}$
In practice, we can also apply annealed Langevin dynamics [4] to facilitate the sampling of each patch, by perturbing it with a series of noise levels $\left\{{{\sigma}_{t}}\right\}_{t\in[TQ/d]}$. Specifically, we refer chained vanilla Langevin dynamics (Chained-VLD) to Algorithm 1 without noise injection (i.e., ${\sigma}_{t}=0$ for all $t\in[TQ/d]$), and chained annealed Langevin dynamics (Chained-ALD) otherwise. Ideally, if a sampler perfectly generates every patch, combining all patches gives a sample from the original distribution due to the chain rule $P({\mathbf{x}})=\prod_{q\in[d/Q]}P({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$. In Proposition 1 we give a linear reduction from producing samples of dimension $d$ using Chained-LD to learning the distribution of a $Q$-dimensional variable for constant $Q$. The proof of Proposition 1 is deferred to Appendix C.
Proposition 1.
@@
in $\tau(\varepsilon/d)$ iterations for some $\varepsilon>0$. Then, equipped with the sampler algorithm, the Chained-LD algorithm in $\frac{d}{Q}\cdot\tau(\varepsilon/d)$ iterations can achieve
@@
With additional assumptions on the target distribution $P$, we can obtain upper bounds on the iteration complexity of Chained-LD. If the conditional distribution $p:=P\left({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots,{\mathbf{x}}^{(q-1)}\right)$ for all $q\in[d/Q]$ satisfies the assumptions specified in Appendix A of [37], i.e., $\log p$ is $L$-Lipschitz smooth for all ${\mathbf{x}}^{(q)}$ and $m$-strongly concave for ${\mathbf{x}}^{(q)}$ outside an $\ell_{2}$-norm ball of radius $R$ (i.e., $\left\|{{\mathbf{x}}^{(q)}}\right\|>R$), by Theorem 1 of [37] we obtain $\tau(\varepsilon/d)={\mathcal{O}}\left(\exp(32LR^{2})\frac{L^{2}}{m^{2}}\cdot\frac{R}{(\varepsilon/d)^{2}}\log\frac{R}{(\varepsilon/d)^{2}}\right)$. In common settings, the radius $R$ in $Q$–dimensional space scales as $R=\Theta(\sqrt{Q})$. Therefore, for constants $L$ and $m$, the iteration complexity of Chained-LD is a polynomial ${\mathcal{O}}\left(\exp({\mathcal{O}}(LQ))\frac{d^{3}}{\varepsilon^{2}\sqrt{Q}}\log\frac{d^{2}\sqrt{Q}}{\varepsilon^{2}}\right)$ in $d$, which takes its minimum value ${\mathcal{O}}\left(\frac{d^{3}}{\varepsilon^{2}}\log\frac{d^{2}}{\varepsilon^{2}}\right)$ when the patch size $Q$ is a constant not growing with $d$. For comparison, the iteration complexity of vanilla Langevin dynamics could scale exponentially in $d$ when $R=\Theta(\sqrt{d})$ in common settings.
Remark 2.
We highlight that Chained-LD is a sampling algorithm. In Algorithm 1, we assume the sampler has direct access to the conditional score function. Also, the conditional densities used in Chained-LD do not require any extra information compared to the target distribution assumed in vanilla LD, since vanilla LD has access to the joint distribution and Chained-LD has access to conditional distributions that, based on chain rule, have the same information as the joint distribution.
Remark 3.
We note that similar combinations of autoregressive models and denoising diffusion models have been studied in the generative modeling literature, in the context of text generation [38, 39] and time series forecasting [40]. At a high level, the generative modeling literature focuses on the training and implementation of autoregressive diffusion models in time-dependent scenarios, while this work focuses on the theoretical guarantees of Chained Langevin dynamics and its comparison with vanilla Langevin dynamics, motivated by their mode-seeking properties under sub-Gaussian mixture distribution.
@@ 6 Numerical Results
Figure 2: Samples from a mixture of three Gaussian modes generated by vanilla Langevin dynamics (VLD) and chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q=10$. Three axes are $\ell_{2}$ distance from samples to the mean of the three modes. The samples are initialized in mode 0.
In this section, we empirically evaluated the mode-seeking tendencies of vanilla and chained Langevin dynamics. We performed numerical experiments on synthetic Gaussian mixture models and real image datasets including MNIST [41] and Fashion-MNIST [42]. Details on the experiment setup are deferred to Appendix D.
Synthetic Gaussian mixture model: We consider the data distribution $P$ as a mixture of three Gaussian components in dimension $d=100$, where mode 0 defined as $P^{(0)}={\mathcal{N}}(\bm{0}_{d},3{\bm{I}}_{d})$ is the in-between mode with high variance, and mode 1 and mode 2 are respectively defined as $P^{(1)}={\mathcal{N}}(\bm{1}_{d},{\bm{I}}_{d})$ and $P^{(2)}={\mathcal{N}}(-\bm{1}_{d},{\bm{I}}_{d})$. The frequencies of the three modes are 0.2, 0.4 and 0.4, i.e.,
@@
In the synthetic experiments, we give the samplers access to the true score function calculated from the target distribution. As shown in Figure 2, vanilla Langevin dynamics (VLD) cannot find mode 1 or 2 within $10^{6}$ iterations if the sample is initialized in mode 0, while chained vanilla Langevin dynamics (Chained-VLD) with patch size $Q=10$ can find the other two modes in 1000 steps and correctly recover their frequencies as gradually increasing the number of iterations. When the sample is initialized in mode 1, as shown in Figure 5 in Appendix D.1, VLD is also likely to be trapped by the high-variance mode 0 and cannot find mode 2, while Chained-VLD is capable of finding all modes. Additional experiments on samples initialized in mode 2 are presented in Appendix D.1, which also verify the mode-seeking tendencies of vanilla Langevin dynamics. We also investigated the effect of different choices of patch size $Q$ on the performance of Chained-LD. As shown in Figures 7, 8, and 9 in Appendix D.1, the convergence of Chained-LD are insensitive to moderate values of constant $Q\in\{1,4,10\}$; for large $Q=20$, it takes more steps to find the other modes; while for overly large $Q=50$, Chained-LD has mode-seeking tendencies similar to LD.
@@
Image datasets: We also perform experiments on generating samples from image datasets by chained annealed Langevin dynamics (Chained-ALD). We construct the distribution as a mixture of two modes by using the original images from MNIST/Fashion-MNIST training dataset (black background and white digits/objects) as the first mode and constructing the second mode by i.i.d. randomly flipping an image (white background and black digits/objects) with probability 0.5.
Since the target distribution of image datasets is unknown, following from [4], we train an estimator to approximate the score function from training samples. More details are deferred to Appendix D.2. We use Recurrent Neural Network (RNN) architectures to estimate the perturbed conditional score function $\nabla_{{\mathbf{x}}^{(q)}}\log P_{{\sigma}_{t}}({\mathbf{x}}^{(q)}\mid{\mathbf{x}}^{(1)},\cdots{\mathbf{x}}^{(q-1)})$ for Chained-ALD. We note that for a sequence of inputs, the output of RNN from the previous step is fed as input to the current step. Therefore, in the scenario of chained Langevin dynamics, the hidden state of RNN contains information about the previous patches and allows the network to estimate the conditional score function of the next patch. More implementation details are deferred to Appendix D.3.
We numerically compare the performance of annealed Langevin dynamics (ALD) and Chained-ALD with different noise levels. The experimental results are shown in Figures 3 and 4. For ALD with bounded noise levels (i.e., the maximum noise $\sigma_{\max}=1$), we observe that it tends to generate the samples from the same mode as initialization, aligning with our theoretical analysis in Theorem 4 in Appendix B. Then, if we apply significantly larger noise levels (i.e., the maximum noise $\sigma_{\max}=50$ as suggested by Technique 1 in [26]), ALD could generate samples from both modes. On the other hand, Chained-ALD, even with bounded noise levels (i.e., $\sigma_{\max}=1$), is capable of finding both modes. Further experiments are deferred to Appendix D.3.
@@
Figure 3: Samples from a mixture distribution of the original and flipped images from the MNIST dataset generated by annealed Langevin dynamics (ALD) and chained annealed Langevin dynamics (Chained-ALD) with patch size $Q=14$ for different numbers of iterations. The maximum noise level $\sigma_{\max}$ is set to be 1 or 50. The samples are initialized as flipped images from MNIST.
@@
Figure 4: Samples from a mixture distribution of the original and flipped images from the Fashion-MNIST dataset generated by annealed Langevin dynamics (ALD) and chained annealed Langevin dynamics (Chained-ALD) with patch size $Q=14$ for different numbers of iterations. The maximum noise level $\sigma_{\max}$ is set to be 1 or 50. The initialization is original images from Fashion-MNIST.
@@ 7 Conclusion
In this work, we theoretically and numerically studied the mode-seeking properties of Langevin dynamics sampling methods under a multi-modal distribution. We characterized Gaussian and sub-Gaussian mixture models under which vanilla Langevin dynamics are unlikely to find all the components within a sub-exponential number of iterations. To reduce the mode-seeking tendency of vanilla Langevin dynamics, we proposed Chained Langevin Dynamics (Chained-LD) and analyzed its convergence behavior. Studying the connections between Chained-LD and denoising diffusion models will be an interesting topic for future exploration. Our RNN-based implementation of Chained-LD is currently limited to image data generation tasks. An interesting future direction is to extend the application of Chained-LD to other domains such as audio and text data. Another future direction could be to study the convergence of Chained-LD under an imperfect score estimation.
@@ Appendix A Theoretical Analysis on the Mode-Seeking Tendency of Langevin Dynamics
We begin by introducing some well-established lemmas used in our proof. We use the following lemma on the tail bound for multivariate Gaussian random variables.
@@ A.1 Proof of Theorem 1: Langevin Dynamics under Gaussian Mixtures
We start by proving that the initialization of the state ${\mathbf{x}}_{0}$ has a large norm on the null space with high probability in the following proposition. Throughout the proof, the notation $\Omega(d)$ refers to $\Omega(d)\geq cd$, for the constant $c$ defined as
$c=\min\left\{\frac{1}{2}\left(\frac{\nu_{0}^{2}-\nu_{\max}^{2}}{8\nu_{0}^{2}}\right)^{2},\frac{1}{8}\left(\log\left(\frac{\nu_{\max}^{2}}{\nu_{0}^{2}}\right)-\frac{\nu_{\max}^{2}}{2\nu_{0}^{2}}+\frac{\nu_{0}^{2}}{2\nu_{\max}^{2}}\right),\frac{1}{32},\frac{(\nu_{0}^{2}-\nu_{\max}^{2})^{2}}{32\nu_{0}^{2}(\nu_{0}^{2}+\nu_{\max}^{2})}\right\},$ (6)
when $d$ satisfies
$d\geq\max\left\{8\left(\log\left(\frac{\nu_{\max}^{2}}{\nu_{0}^{2}}\right)-\frac{\nu_{\max}^{2}}{2\nu_{0}^{2}}+\frac{\nu_{0}^{2}}{2\nu_{\max}^{2}}\right)^{-1}\log\left(\frac{3\nu_{0}^{3}}{w_{0}\min_{i\in[k]}\nu_{i}^{2}}\right),\frac{8\nu_{0}^{2}(3\nu_{0}^{2}+\nu_{\max}^{2})}{\pi(\nu_{0}^{2}-\nu_{\max}^{2})^{2}}\right\}.$ (7)
@@
A.2 Proof of Corollary 1
By the definition of total variation distance, we have
$\text{TV}(\hat{P}_{t},P)=\sup_{A}|\hat{P}_{t}(A)-P(A)|.$
Specifically, by choosing the event $A$ as $\left\{{{\mathbf{x}}:\forall i\in[k],\,\left\|{{\mathbf{x}}-{\bm{\mu}}_{i}}\right\|^{2}\geq\frac{{\nu}_{0}^{2}+{\nu}_{\max}^{2}}{2}d}\right\}$, from Theorem 1 we know $\hat{P}_{t}(A)\geq 1-T\cdot\exp(-\Omega(d))$. On the other hand, by Lemma 1 we have
$P(A)$ $=\sum_{i=0}^{k}w_{i}P^{(i)}(A)$
$\leq w_{0}+\sum_{i=0}^{k}w_{i}\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{\max}^{2}}\right)d\right)$
$=w_{0}+(1-w_{0})\exp\left(-\left(\frac{{\nu}_{0}^{2}-{\nu}_{\max}^{2}}{8{\nu}_{\max}^{2}}\right)d\right).$
Combining the two bounds, we obtain a lower bound on the total variation distance
$\text{TV}(\hat{P}_{t},P)\geq\hat{P}_{t}(A)-P(A)\geq(1-w_{0})\left(1-\frac{T}{\exp(\Omega(d))}\right).$